OpenStudy (anonymous):

I thought after you distribute the 3, you have to add 18 and 3 but then my teacher said you dont , you just then solve for x...im really confused heres the problem, it would be GREAT if someone can explain everything to me y=-2x+6 3y-x+3=0

jimthompson5910 (jim_thompson5910):

Start with the equation 3y-x+3=0 Now replace every copy of y with -2x+6 this is because y = -2x + 6 is the first equation

jimthompson5910 (jim_thompson5910):

so, 3y-x+3=0 3(-2x+6)-x+3=0 3(-2x) + 3(6) - x + 3 = 0 -6x + 18 - x + 3 = 0 -7x + 21 = 0 I'll let you finish in solving for x

jimthompson5910 (jim_thompson5910):

Tell me what you get

OpenStudy (anonymous):

first you have to change the second equation to make it 3y=-x-3 then u add them

OpenStudy (anonymous):

Oh okay so my answer is (3,0)

jimthompson5910 (jim_thompson5910):

very good, (3,0) is correct